OpenStudy (anonymous):

Tara's cell phone plan costs $39.00 a month, which includes 100 text messages. After she uses all of her text messages, it will cost her $.15 per text message

OpenStudy (anonymous):

i have to write an equation that could be used to determine the total cost of her cell phone bill after her first 100 messages.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let x = number of messages over 100 (eg: x = 3 means she sent 103 messages) so if it costs $39 a month + $0.15 per message over 100 messages, then cost = 39 + 0.15*(number of messages over 100) y = 39 + 0.15x where y is the cost and again x is the number of messages over 100

jimthompson5910 (jim_thompson5910):

this equation is only valid if she sends over 100 messages

OpenStudy (anonymous):

well y=39+.15x would be my equation then, the question asks me that if tara only wants to spend $43 how many txt messages can she send?

jimthompson5910 (jim_thompson5910):

replace y with 43 (since y is the cost)

jimthompson5910 (jim_thompson5910):

to get y=39+.15x 43=39+.15x

jimthompson5910 (jim_thompson5910):

now solve for x

OpenStudy (anonymous):

well i subtracted 39 on both sides and got 4=.15x

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

then what

OpenStudy (anonymous):

well i would divide 4 by .15 but it comes out a decimal so then what?

jimthompson5910 (jim_thompson5910):

4/0.15 = 26.667 so x = 26.667 you then round down to get x = 26

OpenStudy (anonymous):

oh alright

jimthompson5910 (jim_thompson5910):

you round down because if you rounded up, your final bill would be over $43